OpenStudy (anonymous):

Using the given points, determine the slope. (8, 1) and (8, -4) slope = 0 slope = -5 slope is undefined i got -6 1/2

OpenStudy (anonymous):

so im guessing its C

OpenStudy (anonymous):

-4-1/8-8 = -5/0 is equals to -infinity so is undefined

zepdrix (zepdrix):

\[\large slope=\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}=\frac{-4-1}{8-8}\]

OpenStudy (anonymous):

oh yes thanks @AzureKais i forgot tht y2 is first not y1

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) m = (-4 - 1) / (8 - 8) m = -5/0 the slope is undefined....this is a vertical line.